# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ A.2 Helper Lemmas
If $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vz,zy\}$ (or by symmetry $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{wy,vz,zy\}$), let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-3=-3$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{vy,vz,zy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path of $\mathsf{OPT}_{1}^{\mathsf{C2}}$, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma 25 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-4=-3$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
@@
If $|D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})|=2$, let w.l.o.g. by symmetry $D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})=\{uy,vy\}$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma 25. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma 24 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma 25 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$. ∎
@@
If the condition in Algorithm 2 in Algorithm 2 applies and $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C3}}$, then there exist edges $F^{\mathsf{C3}}$ with $|F^{\mathsf{C3}}|\leq 4$ such that $H^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|\leq 0$.
@@
First note that the claimed $\mathsf{C3}$ solution in Algorithm 2 is guaranteed by Lemma 24 (among renaming $u$, $v$ and $w$). Let $C_{2}^{\mathsf{C3}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C3}}$ that contains $x$, for $x\in\{u,v,w\}$. We have 4 main cases.
Case 1: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,vw\}|=\ell\leq 2$. First observe that $H_{2}^{\mathsf{C3}}$ is connected as otherwise $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ cannot be feasible for $G_{2}^{\mathsf{C3}}$. Moreover, we can select $\ell$ edges $F^{\mathsf{C3}}\subseteq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ such that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is connected. Further note that every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). Thus, Lemma 25 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=\ell-\ell=0$.
@@
If $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv}\}$ is connected. Every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ between $u$ and $v$, which are also connected in $H_{2}^{\mathsf{C3}}$. Note that there cannot be an edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ on a path between $w$ and $u$ or $v$ that is not in a 2EC component by the choice of the $\mathsf{C3}$ solution $\mathsf{OPT}_{1}^{\mathsf{C3}}$. Moreover, every edge in $H_{2}^{\mathsf{C3}}$ is part of a 2EC component of $H_{2}^{\mathsf{C3}}$ or lies on a path in $H_{2}^{\mathsf{C3}}$ between $u$ and $v$ or $w$, which are also connected in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C3}}$ on a path between $v$ and $w$ that is not in a 2EC component, because $C_{2}^{\mathsf{C3}}(v)=C_{2}^{\mathsf{C3}}(w)$ is 2EC. Thus, Lemma 25 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=1-2\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$ is not connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, by Lemma 24 there must exist an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C3}}(u)$ and $C_{2}^{\mathsf{C3}}(v)$ such that $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-2=0$.

### v2

@@ A.2.2 Auxiliary Lemmas for Algorithm˜3
If $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vz,zy\}$ (or by symmetry $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{wy,vz,zy\}$), let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma˜36. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-3=-3$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{vy,vz,zy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path of $\mathsf{OPT}_{1}^{\mathsf{C2}}$, where $C_{1}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma˜35 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma˜36 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-4=-3$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C2}}$, then there exist edges $F^{\mathsf{C2}}$ with $|F^{\mathsf{C2}}|\leq 1$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ such that $H^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|\leq-2$.
@@
If $|D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})|=2$, let w.l.o.g. by symmetry $D\cap\mathsf{RED}(G_{2}^{\mathsf{C2}})=\{uy,vy\}$ and $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Note that every edge of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}$ between $u$ and $v$ between which is also a path in $H_{2}^{\mathsf{C2}}$. Further, every edge in $H_{2}^{\mathsf{C2}}$ is part of a 2EC component of $H_{2}^{\mathsf{C2}}$ or lies on a path in $H_{2}^{\mathsf{C2}}$ between $u$ and $v$, and there is a path between $u$ and $v$ in $\mathsf{OPT}_{1}^{\mathsf{C2}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C2}}$ that is on a path between $w$ and $u$ or $v$ but not in a 2EC component, because then $\mathsf{RED}(G_{2}^{\mathsf{C2}})$ cannot be feasible for $G_{2}^{\mathsf{C2}}$ as $wy\notin\mathsf{RED}(G_{2}^{\mathsf{C2}})$. Thus, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup H_{2}^{\mathsf{C2}}$ is a 2EC spanning subgraph of $G$ by Lemma˜36. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=0-2=-2$.
Otherwise, that is, $\mathsf{RED}(G_{2}^{\mathsf{C2}})\cap D=\{uy,vy,wy\}$, let $\mathsf{OPT}_{1}^{\mathsf{C2}}$ be an optimal $\mathsf{C2}$ solution for $G_{1}$ that contains a $C_{1}^{\mathsf{C2}}(u)-C_{1}^{\mathsf{C2}}(v)$ path, where $C_{i}^{\mathsf{C2}}(x)$ denotes the 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}$ that contains $x$, for $x\in\{u,v,w\}$. Lemma˜35 implies that there exists an edge $e$ in $G_{1}$ between $C_{1}^{\mathsf{C2}}(w)$ and the connected component containing $V(G_{1})\setminus V(C_{1}^{\mathsf{C2}}(w))$. Therefore, $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ is connected. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ (resp. $H_{2}^{\mathsf{C2}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$ ($H_{2}^{\mathsf{C2}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C2}}$ ($\mathsf{OPT}_{1}^{\mathsf{C2}}\cup\{e\}$). We conclude using Lemma˜36 that $H_{2}^{\mathsf{C2}}\cup\mathsf{OPT}_{1}^{\mathsf{C2}}\cup F^{\mathsf{C2}}$ with $F^{\mathsf{C2}}\coloneq\{e\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C2}}|-|\mathsf{RED}(G_{2}^{\mathsf{C2}})|+|H_{2}^{\mathsf{C2}}|=1-3=-2$. ∎
@@
If the condition in Algorithm˜3 in Algorithm˜3 applies and $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ is a 2EC spanning subgraph of $G_{2}^{\mathsf{C3}}$, then there exist edges $F^{\mathsf{C3}}$ with $|F^{\mathsf{C3}}|\leq 4$ such that $H^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$ and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|\leq 0$.
@@
First note that the claimed $\mathsf{C3}$ solution in Algorithm˜3 is guaranteed by Lemma˜35 (among renaming $u$, $v$ and $w$). Let $C_{2}^{\mathsf{C3}}(x)$ be the 2EC component of $H_{2}^{\mathsf{C3}}$ that contains $x$, for $x\in\{u,v,w\}$. We have 4 main cases.
Case 1: $|\mathsf{RED}(G_{2}^{\mathsf{C3}})\cap\{uv,vw\}|=\ell\leq 2$. First observe that $H_{2}^{\mathsf{C3}}$ is connected as otherwise $\mathsf{RED}(G_{2}^{\mathsf{C3}})$ cannot be feasible for $G_{2}^{\mathsf{C3}}$. Moreover, we can select $\ell$ edges $F^{\mathsf{C3}}\subseteq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ such that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is connected. Further, note that every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). Thus, Lemma˜36 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=\ell-\ell=0$.
@@
If $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv}\}$ is connected. Every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ between $u$ and $v$, which are also connected in $H_{2}^{\mathsf{C3}}$. Note that there cannot be an edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ on a path between $w$ and $u$ or $v$ that is not in a 2EC component by the choice of the $\mathsf{C3}$ solution $\mathsf{OPT}_{1}^{\mathsf{C3}}$. Moreover, every edge in $H_{2}^{\mathsf{C3}}$ is part of a 2EC component of $H_{2}^{\mathsf{C3}}$ or lies on a path in $H_{2}^{\mathsf{C3}}$ between $u$ and $v$ or $w$, which are also connected in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$. Note that there cannot be an edge in $H_{2}^{\mathsf{C3}}$ on a path between $v$ and $w$ that is not in a 2EC component, because $C_{2}^{\mathsf{C3}}(v)=C_{2}^{\mathsf{C3}}(w)$ is 2EC. Thus, Lemma˜36 implies that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=1-2\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$ is not connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$, by Lemma˜35 there must exist an edge $e$ in $G_{2}$ between $C_{2}^{\mathsf{C3}}(u)$ and $C_{2}^{\mathsf{C3}}(v)$ such that $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Thus, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-2=0$.